OpenStudy (anonymous):

Solve the absolute value equation. Check your solutions. | 5x + 13| = –7

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

There will be two separate equations to solve from: First when the 5x+13 is positive which means x>-13/5 which leads to +(5x+13)=-7 solving it gives the value as -4 since we 've assumed the above eqn for case x>-13/5 thus -4 is discarded. Now another eqn comes when 5x+13 is -ve means x<-13/5 which leads to -(5x+13)=-7 which gives us a solution as x=-6/5 which is discarded because previously we 'd assumed it for x<-13/5 but -6/5 is greater than -13/5 Hence there is no solution to this equation. We can prove it by graphically also. .